Phosgene is a potent chemical warfare agent that is now outlawed by international agreement. It decomposes by the reaction: COCl2 (g) ⇋   CO (g) + Cl2 (g) Kc = 4.1 x 10-5 (at 359°C)

Calculate the concentration of CO when 7.45 mol of phosgene decomposes and reaches equilibrium in a 10.0 liter flask. Multiply your answer by 103 and enter that number to 2 decimal places.
Answer #1

Moles of CO = 7.45 mol

Volume of Flask = 10L

Molarity of COCl2 = Moles/Volume = 7.45 mol / 10 L = 0.745 M

Given reaction is:

COCl2 (g) ⇋   CO (g) + Cl2 (g)

I(M) 0.745 0 0

C(M) -x +x +x

E(M) 0.745-x x x

Kc = [CO][Cl2] /[COCl2]

4.1 *10-5 = x2 / (0.745-x)

4.1 *10-5 (0.745 -x) = x2

x2 + (4.1 *10-5)x - (3.05 *10-5) = 0

x = 0.0055 = 5.5 *10-3

Thus, at equilibrium , [CO] = x = 5.5 *10-3 M

Now, if we multiply by 103, then answer is 5.5